OpenStudy (sasogeek):

\(\huge 3^x - 2^{y+2}=49 \) \(\huge 2^y - 3^{x-2}+1=0 \) solve the system of equations.

OpenStudy (sasogeek):

i tried using logs but i got to a dead end... :/

OpenStudy (asnaseer):

try using a subtle substitution

OpenStudy (sasogeek):

how?

OpenStudy (asnaseer):

\[a=3^x\]\[b=2^y\]

OpenStudy (asnaseer):

remember that \(2^{y+2}=2^2\times2^y=4\times2^y\)

OpenStudy (sasogeek):

ohhh, ok :) yeah...

OpenStudy (asnaseer):

:)

OpenStudy (turingtest):

This is why I love this site, I never would have known what to do there...

OpenStudy (sasogeek):

i know right, i'm helping a friend out but i'm stuck myself so thanks to you guys ;)

OpenStudy (asnaseer):

I agree - I have learnt SO MUCH from this site - it is truly amazing.

OpenStudy (anonymous):

where's the button for "round of applause"?

OpenStudy (sasogeek):

lol

OpenStudy (asnaseer):

he he :D

OpenStudy (sasogeek):

thanks guys, we did it! xD \(\large x \approx 4.069 \) and \(\large y \approx 3.263\) can someone confirm this please... :)

OpenStudy (asnaseer):

hmmm - I actually get integer solutions

OpenStudy (asnaseer):

Using the substitutions I listed above, I get:\[a-4b=49\tag{1}\]\[b-\frac{a}{9}+1=0\tag{2}\]Multiplying (2) by 9 and moving the constant to the right-hand-side, we get:\[9b-a=-9\tag{3}\]Adding (1) and (3) gives:\[5b=40\implies b=8\]Substituting this into (1) gives:\[a=81\]Therefore:\[2^y=8\implies y=3\]\[3^x=81\implies x=4\]